OpenStudy (anonymous):

9 ( k - 2 ) = 7 ( k - 8 ) solving proportions

OpenStudy (anonymous):

9k-18=7k-56 2k=-38 k=-19

OpenStudy (anonymous):

9k-18=7k-56 2k= you are a superhero! you are fast

OpenStudy (anonymous):

so k = -19

OpenStudy (anonymous):

yes